OpenStudy (anonymous):

g(2) = 6 and f(6) = 18, then (f o g)(2)

OpenStudy (gorv):

(fog)(2)=f(g(2)) g(2)=6

OpenStudy (gorv):

f(6)=???

OpenStudy (gorv):

we plugged in the value of g(2) in f(g(2))

OpenStudy (gorv):

g(2)=6 so replace g(2) by 6

OpenStudy (gorv):

f(g(2))=f(6)=?

OpenStudy (mathmath333):

\(\textrm{here is one method}\) \(\large\tt \begin{align} \color{black}{g(2) = 6 \\ \text{hence}\\ g(x)=3x\\ f(6) = 18\\ \text{hence}\\ f(x)=3x\\ (f o g)(x)=3(3x)\\ (f o g)(2)=3(3\times 2)\\ =18\\}\end{align}\)